Solve the equation x+y+z=6, x+2y+3z=14, x+4y+9z=36
Answer by TimothyLamb(4379)

x+y+z=6, x+2y+3z=14, x+4y+9z=36
---
put the system of linear equations into standard form
---
x+y+z=6
x+2y+3z=14
x+4y+9z=36
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 1
y = 2
z = 3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php